A basketball floats in a bathtub of water. The ball has a mass of 0.5 kg and a diameter of 22 cm.




(a) What is the buoyant force?




(b) What is the volume of water displaced by the ball?




(c) What is the average density of the basketball?




You can use g = 9.8 m/s2




.





1
Expert's answer
2022-04-06T18:36:47-0400

"F=\rho_w g\frac{\pi d^3}6=54.6~N,"

"V=\frac m{\rho_w}=0.5~dm^3,"

"\rho_b=\frac{6m}{\pi d^3}=9~\frac{kg}{m^3}."